A 100kg sled starts from rest on an icy surface. A team of dogs accelerates the sled uniformly to 5.0m/s over a period of 5.0s. (a)[4 pt(s) ]What is the force supplied by the dogs on the sled? (b)[5 pt(s) ]What is the amount of work done by the dogs on the sled in that 5.0s time? (c)[2 pt(s) ]What does impulse measure? (d)[4 pt(s) ]What is the impulse delivered to the sled by the dogs over the entire time t = 5.0s? (e)[4 pt(s) ]What is the instantaneous power due to the dogs at time t = 2.5s?
Answer #1

The acceleration is

a = (5-0)/5 = 1m/s

The force supplied by the dogs is

F = m a = (100)(1) = 100 N

Distance travelled in 5seconds is

d = (1/2)at^2 = (1/2)(1)(5)^2 = 12.5m

Work done is W = F. d = (100)(12.5) = 1250 J

Impulse measures the change in momentum

I = m (vf-vi) = 100(5-0) = 500 kgm/s

the impulse delivered to the sled by the dogs over the entire time

IMpulse I = F. t = (100)(5) =500 Ns

instantaneous power due to the dogs at time

P = W/ t = 1250/2.5 = 500 w